40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 29 22:37:54 2022, Koji, Update, IOO, WFS investigation PXL_20220730_040900843.jpgPXL_20220730_041216848.jpg
    Reply  Tue Aug 2 01:10:26 2022, Koji, Update, IOO, WFS investigation 220801_IMC_WFS_DEMOD.pdf220801_IMC_WFS_DEMOD2.pdfWFS1.pngWFS2.png
       Reply  Thu Aug 4 21:58:18 2022, Koji, Update, IOO, WFS investigation Screenshot_2022-08-04_22-10-57.pngScreenshot_2022-08-04_22-11-16.pngScreenshot_2022-08-04_22-11-53.pngScreenshot_2022-08-04_22-12-39.png
          Reply  Thu Aug 4 22:14:20 2022, Koji, Update, IOO, WFS investigation IMC_SUS_channels_TF.pdfIMC_WFS_channels_TF.pdfIMC_WFS_segment_TF.pdfIMC_WFS_220804.xlsx
             Reply  Thu Aug 4 22:14:38 2022, Koji, Update, IOO, WFS investigation Screenshot_2022-08-04_22-17-19.pngScreenshot_2022-08-04_22-18-45.png
                Reply  Thu Aug 4 22:32:31 2022, Koji, Update, IOO, Upon leaving the lab (WFS investigation) 
                   Reply  Fri Aug 5 12:42:12 2022, Koji, Update, IOO, IMC WFS: Overnight observation  Screen_Shot_2022-08-05_at_12.04.01.png
Message ID: 17061     Entry time: Thu Aug 4 22:14:38 2022     In reply to: 17060     Reply to this: 17062
Author: Koji 
Type: Update 
Category: IOO 
Subject: WFS investigation 

WFS/MCTRANS_QPD Power Spectra

Attachment 1: HEPA ON

WFS1/2 PIT/YAW Spectra are stabilized below 1Hz (0.1Hz for WFS1P)

MC2 TRANS PIT is largely contaminated by the other WFS loops.
MC2 TRANS YAW is slightly contaminated but not much compared to the one for pitch.

Attachment 2: HEPA OFF

Again, WFS1/2 PIT/YAW Spectra are stabilized below 1Hz (0.1Hz for WFS1P)

MC2 TRANS PIT is still contaminated but better.
MC2 TRANS YAW is not contaminated.


Observation

WFS1/2 signals are largely disturbed when PSL HEPA is ON. Probably the amount of HEPA air flow was not optimized.
Above 1Hz, invacuum suspension are quieter than the beam incident on the IMC.

The dirty WFS signals are fedback to the mirrors. Due the large motion of the beam and also the imperfection of the actuator matrix cause the MC2 spot rather moves than stabilized.

This means that the WFS loops should leave the mirrors untouched above 1Hz i.e. The loop bandwidths should be low (~<0.1Hz). (Yes I know)
However, the simple gain reduction (x10) does make the servos unstable. So more adjustment is necessary. (<-Not for today)

Attachment 1: Screenshot_2022-08-04_22-17-19.png  306 kB  Uploaded Fri Aug 5 01:58:08 2022  | Hide | Hide all
Screenshot_2022-08-04_22-17-19.png
Attachment 2: Screenshot_2022-08-04_22-18-45.png  305 kB  Uploaded Fri Aug 5 01:58:12 2022  | Hide | Hide all
Screenshot_2022-08-04_22-18-45.png
ELOG V3.1.3-